OpenStudy (anonymous):

f(x)= (x^3-1) ln |x|.find f' (x)

OpenStudy (anonymous):

product rule for this one

OpenStudy (anonymous):

\((fg)'=f'g+g'f\) with \[f(x)=x^3-1,f'(x)=3x^2,g(x)=\ln(x),g'(x)=\frac{1}{x}\]

OpenStudy (anonymous):

Thank you. I have another question to ask. Can you help me, please? That would be the last one.

OpenStudy (anonymous):

4. Find the value of c such that ∫_0^c e^-2x dx=1/4